OpenStudy (eamier):

S:x^2+y^2+z^2-2x-4y-4x=0, π:2x+y+2z-5=0 find the circle area of sphere S cut by plane π.

OpenStudy (turingtest):

-2x-4y-4x=-4y-6x sure that's not a typo? why is it unsimplified?

OpenStudy (turingtest):

-2x-4y-4z maybe?

OpenStudy (eamier):

i don't know the answer. not typo

OpenStudy (eamier):

the answer must be circle i think.

OpenStudy (turingtest):

S:x^2+y^2+z^2-2x-4y-4x=0 ^ ^ are you *sure* these are supposed to both be x's ?

OpenStudy (turingtest):

it's just weird to have a problem like that, where the first step is algebra I level simplification...

OpenStudy (eamier):

S is sphere, and pi is plane. when the plane cut the the sphere we get a circle. is it true to think like this

OpenStudy (turingtest):

ok well if you insist that there is no typo, yes it is a sphere we can complete the square in x and y to get it into standard form

OpenStudy (turingtest):

hm... I have to think about getting the intersect a bit...

OpenStudy (eamier):

I graphed the equations.. still dunno how

OpenStudy (turingtest):

we gotta eliminate one of the variables somehow with a substitution from the plane equation into the sphere one

OpenStudy (turingtest):

hey that is a z... you said there were no typos :/

OpenStudy (turingtest):

you wrote x^2+y^2+z^2-2x-4y-4x above and x^2+y^2+z^2-2x-4y-4z on your link

OpenStudy (turingtest):

ok, since we have\[2x+y+2z-5=0\]we also have that\[10-4x-2y-4z=0\](multiplication by -2) since this and the sphere equation are both equal to zero we can set them equal to each other\[10-4x-2y-\cancel{4z}=x^2+y^2+z^2-2x-4y-\cancel{4z}\]we can cancel a bit\[x^2+y^2+z^2+2x-2y-10=0\]now solve for z in the plane equation and substitute that in...

OpenStudy (turingtest):

\[z=\frac52-x-\frac12y\]\[x^2+y^2+(\frac52-x-y)^2+2x-2y-10=0\]hm... now I am thinking that we won't be able to complete the square, which was what I wanted to do here...

OpenStudy (turingtest):

ok I just read an interesting approach from wikipedia let me try it out! http://en.wikipedia.org/wiki/Plane%E2%80%93sphere_intersection

OpenStudy (eamier):

ok ill try too

OpenStudy (turingtest):

first we need the radius of the sphere, which we can find by completing the square (or with software of course)

OpenStudy (turingtest):

do you want to see how to get it by completing the square?

OpenStudy (turingtest):

I'll just imagine that you know it I guess... then you need the equation of the plane in the form\[ax+by+cz=d\]such that\[a^2+b^2+c^2=1\]since for yours we get\[2x+y+2z=5\]we need to change the coefficients, which we do my dividing by the magnitude of the gradient of the plane

OpenStudy (turingtest):

\[\|\nabla f\|=\sqrt{2^2+1^2+2^2}=3\]so we get\[\frac23x+\frac13y+\frac23z=\frac53=d\]and the radius of the sphere is \[r=3\]

OpenStudy (turingtest):

according to the article, it says of the intersection of these forms "if d < r, (it forms) a circle of intersection, with center (da, db, dc) and squared radius r^2 − d^2."

OpenStudy (turingtest):

so apparently the square of the radius of that circle that is formed by the intersection should be\[R^2=r^2-d^2=3^2-(\frac53)^2=\frac{81}9-\frac{25}9=\frac{56}9\]hence \[A=\pi R^2=\frac{56\pi}9\]

OpenStudy (turingtest):

at lest that's what I get from the wiki article...

OpenStudy (turingtest):

least*

OpenStudy (eamier):

you really help a lot.

OpenStudy (turingtest):

thanks, I really learned a lot :)

OpenStudy (anonymous):

*

OpenStudy (eamier):

what's that mean